# Leveraging Reinforcement Learning, Genetic Algorithms and Transformers for background determination in particle physics

### v1

@@ Abstract
This paper has two primary goals. First, from a particle physics perspective, we present a novel approach that utilises Reinforcement Learning (RL) to overcome the aforementioned challenges by systematically determining the critical backgrounds affecting beauty hadron decay measurements. While beauty hadron physics serves as the case study in this work, the proposed strategy is broadly adaptable to other types of particle physics measurements. Second, from a Machine Learning perspective, we introduce a novel algorithm which exploits the synergy between RL and Genetic Algorithms (GAs) for environments with highly sparse rewards and a large trajectory space. This strategy leverages GAs to efficiently explore the trajectory space and identify successful trajectories, which are used to guide the RL agent’s training. Our method also incorporates a transformer architecture for the RL agent to handle token sequences representing decays.
@@ 1 Introduction
In particle physics, measurements are often framed in terms of a signal and its accompanying backgrounds. The signal corresponds to the specific process under study, such as the production of a new particle or a rare interaction, while backgrounds are all other processes that can mimic the same experimental signature but are not of direct interest. An important area is heavy flavor physics, where beauty hadron decays serve as a case study for precision measurements and searches for rare processes to test the Standard Model and probe for signs of new physics, with experiments such as LHCb and B-factories like Belle II playing a leading role. Beauty hadrons can decay into a large number of final states, often through cascades involving several intermediate particles, and given the vast number of possible decay channels and their overlapping kinematic signatures, identifying the most relevant backgrounds for a given signal is both critical and highly non-trivial.
@@ 2 Related works
Recent years have seen a growing interest in leveraging advanced Artificial Intelligence (AI) methods to address complex problems in particle physics, with RL, transformers, and GAs emerging as prominent tools. RL has found important applications in the field of particle physics. Examples include instrument design [2], particle tracking [3], and particle physics model building [4, 5]. In our prior work, we introduced the DL Advocate framework [6], which applies RL to quantitatively assess the impact of hidden systematic uncertainties in precision measurements. Concurrently, the field of RL has witnessed significant algorithmic developments. A particularly relevant advancement is the AlphaZero algorithm [7], which outperformed world champion programs in the games of chess, shogi and Go. This algorithm starts the learning process with no domain knowledge except the game rules, and employs self-play to generate the training data. It combines deep neural networks with Monte Carlo Tree Search (MCTS) to evaluate game positions and possible moves. This strategy facilitates the discovery of promising action trajectories, which is particularly valuable in environments with large trajectory spaces and delayed rewards, where feedback is only obtained at the end of an episode, as in the problem addressed in this work. While AlphaZero has seen limited use in physics so far, recent works have begun to explore its application, for instance, in symbolic regression to find analytical methods within the field [8], and in the optimisation of quantum dynamics [9]. However, to the best of our knowledge, the present work constitutes the first application of the AlphaZero algorithm within the domain of particle physics.
@@
Additionally, we studied an alternative to PGSU for Phase 2, which we named Periodic Expert Guidance (PEG), consisting of adjusting the action masking so that the agent is forced to follow expert trajectories during certain episodes of training. Employing this strategy only for a fraction of episodes allows the agent to explore beyond these trajectories and achieve beyond-expert performance333In this work, performance is measured as the number of backgrounds correctly determined.. Although PGSU typically achieved higher final performance in the environment of this work, PEG can yield comparable results while providing faster initial performance improvement, as we only added a supervised term to the loss in Equation 1 for states along expert trajectories. This requires the agent to learn how to reach a state before it can receive supervision for that state. Consequently, PEG could be particularly useful in environments with very long episodes, where PGSU might lead to a slower learning.
@@ 3.1.3 Action masking
Without action masking, the size of the trajectory space in the described RL problem is upper-bounded by , where denotes the number of actions and the maximum sequence length allowed to represent a background. This results in an intractably large space of possible trajectories, except in extremely simplified scenarios. However, by employing action masking, the effective trajectory space can be substantially reduced. This is possible due to the structured nature of the tokenization scheme and constraints imposed by physical laws, which allow for the elimination of invalid or redundant actions. For instance, when generating an intermediate resonance, the closing parenthesis token is masked unless charge, lepton flavour, and other conservation laws are satisfied. Similarly, the “End” token is masked unless the number of intermediate particle tokens in the sequence matches the number of closing parenthesis, ensuring syntactic consistency. Action masking not only enhances computational efficiency but also improves the agent’s exploration and learning capabilities.
@@ 3.1.4 RL algorithm
The RL problem described presents several key challenges: despite the application of action masking, the trajectory space remains large, contributing to the complexity of exploration. Additionally, the reward is only provided at the end of the episode, complicating the learning process of the agent, as it receives no immediate feedback when selecting an action. Moreover, the reward structure is highly sparse: only a small fraction of sequences correspond to valid background processes, making successful trajectories rare. In this context, the RL algorithm that we chose is AlphaZero [7], which has demonstrated its effectiveness in games like chess, shogi and Go, all of which involve large trajectory spaces and delayed rewards. The algorithm is originally designed for two-player adversarial games. In this work, however, we adapt an open-source AlphaZero implementation [29] to operate in a single-player setting, specifically for solving the background finder problem.
@@ 3.1.5 Synergy of RL and GAs
In equation 3, indexes an expert trajectory, is the number of expert trajectories for the corresponding training signal, the number of timesteps for which the agent followed trajectory , the reward associated with trajectory , a timestep, the expert action at timestep , and the corresponding state. The negative log-likelihood penalises the model whenever it assigns low probability to expert actions. This supervised loss is applied only to states along expert trajectories, and the loss of each trajectory is normalised by the number of states for which the agent followed the trajectory, rather than the full trajectory length. This ensures that the loss decreases as training progresses: without this normalisation, the loss could initially increase as the agent begins to follow longer portions of expert trajectories, since more terms would be added to the loss. Finally, it is important to note that GAs may not discover all relevant backgrounds for the training signals. For this reason, we adopt the PGSU approach rather than BC, as the former enables exploration beyond expert demonstrations and allows achieving beyond-expert performance.
@@ 3.1.6 Determination of the relevant backgrounds
Once the training of the agent is complete, the backgrounds corresponding to a new signal are determined by sampling actions from the probability distribution provided by the model until the “End” token is selected. This sampling process is repeated multiple times to obtain a set of token sequences. If training has been effective and the agent has achieved good generalisation, actions leading to a relevant background sequence will be assigned high probabilities, and the decay will likely be part of the sampled set. To promote diversity, a high sampling temperature is used (see Equation 4, where is the original probability for action , and is the adjusted probability after applying temperature scaling), which encourages the selection of actions beyond those with the highest probability, and thus target the determination of all relevant backgrounds.
@@ 3.1.8 Model fine tuning
After completing the training phase, we implemented a fine-tuning stage to further improve the model’s performance. During training, the algorithm records the relevant background configurations encountered by the agent. In the fine-tuning stage, the agent employs these trajectories to specialise on this knowledge. As in training, we designed two distinct strategies: for the agent trained with PGSU, fine tunining consists of performing a second PGSU employing all relevant background sequences built during training as expert demonstrations (rather than only the GA-discovered trajectories). In the fine tuning of the agent trained with PEG, the agent is required to play multiple episodes focusing solely on these specific configurations.
@@ 3.2.2 Offspring generation and natural selection
The next generation is obtained through natural selection. First, elite individuals666The elite refers to the best individuals in the current population of a specific generation, typically preserved for the next generation to ensure that high-quality solutions are not lost. On the other hand, the hall of fame is a mechanism to track and preserve the best individuals across multiple generations, not just within a single generation. are identified. These are individuals with fitness values sufficiently high to be considered relevant backgrounds, and they are directly carried over to the next generation. Additionally, tournament selection is employed within the population. In each tournament, three individuals are randomly chosen, and the individual with the highest fitness value is selected to advance to the next generation. This selection continues until the new generation matches the size of the previous population.
@@ 3.2.4.1 Intermediate resonance construction
We designed two different strategies for selecting the intermediate particle: a naive approach, in which the particle is randomly sampled from the set of intermediate particles with the appropriate charge, and a more sophisticated method that relies on decay channels already incorporated into the algorithm’s internal knowledge base. The designed GAs collect knowledge about existing decays in a dictionary structure. Each key in the dictionary corresponds to a particle, and the associated value is the set of allowed decay channels for that particle. This dictionary is initialised with decay information from the PDG. However, if a decay not already present in the dictionary is encountered during optimisation, the GA learns777Not to be confused with the meaning of learning in ML. and incorporates this new information. In contrast to conventional GA applications, where problems are typically independent by nature (the optimisation of a function does not affect the optimisation of a different function ), this property of independence does not hold in the context of the background finder: the GA may identify a background involving a decay not listed in the PDG, and this newly discovered information could be relevant for other signals beyond the one currently under consideration. After the optimisation for a given signal is completed, the dictionary can be saved and reused for background identification in subsequent signals, avoiding redundant rediscovery of decays.
@@ 4 Environment
The misidentification penalty, , is defined as , where denotes the number of misidentified particles.
@@ 5.1 GA performance
We conducted an initial experiment to evaluate the effectiveness of Genetic Algorithms (GAs) to identify the backgrounds associated with the training signals. Based on the toy model reward function, 94 backgrounds are considered relevant. To assess GA performance, defined as the number of relevant backgrounds identified, we performed multiple runs using varying population sizes and numbers of generations. The results, presented in Figure 8, demonstrate a clear upward trend in performance as either hyperparameter increases888Background processes may contain more final state particles than the signal if they involve non-reconstructed particles. For a more efficient search, instead of allowing the number of positive, negative and neutral final state particles in individuals to vary freely, we fixed these counts for the entire population. As a consequence, the process of finding the relevant backgrounds for a signal will in reality involve running multiple GA evolutions, each with a different configuration of these counts. For the signals in the considered training set, the number of evolutions is 3 or 4 depending on the case. The reported population sizes refer to those used in each of these evolutions.. This trend is expected: larger populations enhance genetic diversity, while more generations allow the algorithm additional opportunities to explore the space. In Appendix B, we compute the GA space size for an example signal to be 333,305,600. The results also indicate that, even with a total number of evaluations much smaller than the GA space size, the algorithm successfully identifies the majority of backgrounds relevant to the training signals.
@@ 5.4 Visualisation of learnt embeddings
6 Conclusion
@@
We developed a toy scenario with 40 particles, using B-mesons as mother particles and experimentally well-established intermediate states, to provide realistic and sufficiently complex decay chains for evaluating the performance of the proposed algorithm. We conducted a first experiment to assess the performance of the GA, where results demonstrate that the GA can successfully identify most of the backgrounds for the training signals, despite using a much smaller number of evaluations than the size of the GA search space.
By employing the GA-discovered backgrounds as expert demonstrations, we trained and fine-tuned two agents, one following the PGSU approach and the other using PEG. The results indicate that both methods surpass expert perfomance and achieve perfect training performance999In this context, perfect performance corresponds to correctly determining every background.. PGSU exhibits robust performance with slightly superior generalisation ability. In contrast, PEG demonstrates a faster initial performance improvement, making it an interesting alternative for environments involving very large trajectories, where PGSU may result in a slower training.

### v2

@@ Abstract
This paper has two primary goals. First, from a particle physics perspective, we present a novel approach that utilises Reinforcement Learning (RL) to overcome the aforementioned challenges by systematically determining the critical backgrounds affecting beauty hadron decay measurements. While beauty hadron physics serves as the case study in this work, the proposed strategy is broadly adaptable to other types of particle physics measurements. Second, from a Machine Learning perspective, we introduce a novel algorithm which exploits the synergy between RL and Genetic Algorithms (GAs) for environments with highly sparse rewards and a large trajectory space. This strategy leverages GAs to efficiently explore the trajectory space and identify successful trajectories, which are used to guide the RL agent’s training. Our method also incorporates a transformer architecture for the RL agent to process token sequences that represent particle decays.
@@ 1 Introduction
In particle physics, measurements are often framed in terms of a signal and its accompanying backgrounds. The signal corresponds to the specific process under study, such as the production of a new particle or a rare interaction, while backgrounds are all other processes that can mimic the same experimental signature but are not of direct interest. An important area is heavy flavor physics, where beauty hadron decays serve as a case study for precision measurements and searches for rare processes to test the Standard Model and probe for signs of new physics, with experiments such as LHCb [1] and B-factories like Belle II [2] playing a leading role. Beauty hadrons can decay into a large number of final states, often through cascades involving several intermediate particles, and given the vast number of possible decay channels and their overlapping kinematic signatures, identifying the most relevant backgrounds for a given signal process under study is both critical and highly non-trivial.
@@ 2 Related works
Recent years have seen a growing interest in leveraging advanced Artificial Intelligence (AI) methods to address complex problems in particle physics, with RL, transformers, and GAs emerging as prominent tools. RL has found important applications in the field of particle physics. Examples include instrument design [4], particle tracking [5], and particle physics model building [6, 7, 8, 9]. In our prior work, we introduced the DL Advocate framework [10], which applies RL to quantitatively assess the impact of hidden systematic uncertainties in precision measurements. Concurrently, the field of RL has witnessed significant algorithmic developments. A particularly relevant advancement is the AlphaZero algorithm [11], which outperformed world champion programs in the games of chess, shogi and Go. This algorithm starts the learning process with no domain knowledge except the game rules, and employs self-play to generate the training data. It combines deep neural networks with Monte Carlo Tree Search (MCTS) [12, 13, 14] to evaluate game positions and possible moves. This strategy facilitates the discovery of promising action trajectories, which is particularly valuable in environments with large trajectory spaces and delayed rewards, where feedback is only obtained at the end of an episode, as in the problem addressed in this work. While AlphaZero has seen limited use in physics so far, recent works have begun to explore its application, for instance, in symbolic regression to find analytical methods within the field [15], and in the optimisation of quantum dynamics [16]. However, to the best of our knowledge, the present work constitutes the first application of the AlphaZero algorithm within the domain of particle physics.
@@
Additionally, we studied an alternative to PGSU for Phase 2, which we named Periodic Expert Guidance (PEG), consisting of adjusting the action masking so that the agent is forced to follow expert trajectories during certain episodes of training. Employing this strategy only for a fraction of episodes allows the agent to explore beyond these trajectories and achieve beyond-expert performance333In this work, performance is measured as the number of backgrounds correctly determined..
@@ 3.1.3 Action masking
Without action masking, the size of the trajectory space in the described RL problem is upper-bounded by , where denotes the number of actions and the maximum sequence length allowed to represent a background. This results in an intractably large space of possible trajectories, except in extremely simplified scenarios. However, by employing action masking, which filters out invalid actions at each decision step, the effective trajectory space can be substantially reduced. This is possible due to the structured nature of the tokenization scheme and constraints imposed by physical laws, which allow for the elimination of invalid or redundant actions. For instance, when generating an intermediate resonance, the closing parenthesis token is masked unless charge, lepton flavour, and other conservation laws are satisfied. Similarly, the “End” token is masked unless the number of intermediate particle tokens in the sequence matches the number of closing parenthesis, ensuring syntactic consistency. Action masking not only enhances computational efficiency but also improves the agent’s exploration and learning capabilities.
@@ 3.1.4 RL algorithm
The RL problem described presents several key challenges: despite the application of action masking, the trajectory space remains large, contributing to the complexity of exploration. Additionally, the reward is only provided at the end of the episode, complicating the learning process of the agent, as it receives no immediate feedback when selecting an action. Moreover, the reward structure is highly sparse: only a small fraction of sequences correspond to valid background processes, making successful trajectories rare. In this context, the RL algorithm that we chose is AlphaZero [11], which has demonstrated its effectiveness in games like chess, shogi and Go, all of which involve large trajectory spaces and delayed rewards. The algorithm is originally designed for two-player adversarial games. In this work, however, we adapt an open-source AlphaZero implementation [36] to operate in a single-player setting, specifically for solving the background finder problem. The traditional AlphaZero loss, defined over the entire set of states in the self-play dataset , is given by Equation 3.
Here, denotes a state, the network’s prediction of the expected outcome for state , the episode outcome, the model’s predicted action probabilities for state , and the policy obtained from MCTS. The value loss measures how accurately the model predicts the final episode outcome, while the policy loss encourages the network’s policy head to match the policy obtained from MCTS. Additionally, the term denotes an regularization term applied to the network parameters, which prevents overfitting and stabilises training.
@@ 3.1.5 Synergy of RL and GAs
In Equation 4, indexes an expert trajectory, and denotes the number of expert trajectories associated with the trajectory , which depends on the corresponding training signal. is the number of timesteps for which the agent followed the expert trajectory , the reward associated with trajectory , a timestep, the policy head’s predicted probability (under parameters ), the expert action at timestep for trajectory , and the corresponding state. The negative log-likelihood penalises the model whenever it assigns low probability to expert actions. This supervised loss is applied only to states along expert trajectories, and the loss of each trajectory is normalised by the number of states for which the agent followed the trajectory, rather than the full trajectory length. This ensures that the loss decreases as training progresses: without this normalisation, the loss could initially increase as the agent begins to follow longer portions of expert trajectories, since more terms would be added to the loss. Finally, it is important to note that GAs may not discover all relevant backgrounds for the training signals. For this reason, we adopt the PGSU approach rather than BC, as the former enables exploration beyond expert demonstrations and allows achieving beyond-expert performance.
The supervised loss in Equation 4 is defined over a single trajectory . Equation 5 specifies the total loss, which corresponds to the sum of the traditional AlphaZero loss [11] and the supervised term, and extends this definition to the entire set of trajectories in the self-play dataset . The expectation in the AlphaZero term averages over all states encountered during self-play, whereas the expectation in the supervised term averages over all self-play trajectories.
Given that training is performed on batches of individual states rather than full trajectories, it is more convenient to express the supervised loss as an expectation over states. In Equation 6, denotes the number of expert trajectories associated with the training signal for state , the expert action prescribed by trajectory for state , and if state appears in expert trajectory , and otherwise 666Note that the numerical value of differs between Equations 5 and 6, as the former averages over the number of trajectories in the dataset, whereas the latter averages over the number of states..
@@ 3.1.6 Determination of the relevant backgrounds
Once the training of the agent is complete, the backgrounds corresponding to a new signal are determined by sampling actions from the probability distribution provided by the model until the “End” token is selected. This sampling process is repeated multiple times to obtain a set of token sequences. If training has been effective and the agent has achieved good generalisation, actions leading to a relevant background sequence will be assigned high probabilities, and the decay will likely be part of the sampled set. To promote diversity, a high sampling temperature () is used (see Equation 7, where is the original probability for action , and is the adjusted probability after applying temperature scaling), which encourages the selection of actions beyond those with the highest probability, and thus target the determination of all relevant backgrounds.
@@ 3.1.8 Model fine tuning
After completing the training phase, we implemented a fine-tuning stage to further improve the model’s performance. During training, the algorithm records the relevant background configurations encountered by the agent. In the fine-tuning stage, the agent employs these trajectories to specialise on this knowledge. As in training, we designed two distinct strategies: for the agent trained with PGSU, fine tuning consists of performing a second PGSU employing all relevant background sequences built during training as expert demonstrations (rather than only the GA-discovered trajectories). In the fine tuning of the agent trained with PEG, the agent is required to play multiple episodes focusing solely on these specific configurations.
@@ 3.2.2 Offspring generation and natural selection
The next generation is obtained through natural selection. First, elite individuals777The elite refers to the best individuals in the current population of a specific generation, typically preserved for the next generation to ensure that high-quality solutions are not lost. On the other hand, the hall of fame is a mechanism to track and preserve the best individuals across multiple generations, not just within a single generation. are identified. These are individuals with fitness values sufficiently high to be considered relevant backgrounds, and they are directly carried over to the next generation. Additionally, tournament selection is employed within the population. In each tournament, three individuals are randomly chosen, and the individual with the highest fitness value is selected to advance to the next generation. This selection continues until the new generation matches the size of the previous population.
@@ 3.2.4.1 Intermediate resonance construction
We designed two different strategies for selecting the intermediate particle: a naive approach, in which the particle is randomly sampled from the set of intermediate particles with the appropriate charge, and a more sophisticated method that relies on decay channels already incorporated into the algorithm’s internal knowledge base. The designed GAs collect knowledge about existing decays in a dictionary structure. Each key in the dictionary corresponds to a particle, and the associated value is the set of allowed decay channels for that particle. This dictionary is initialised with decay information from the PDG. However, if a decay not already present in the dictionary is encountered during optimisation, the GA learns888Not to be confused with the meaning of learning in ML. and incorporates this new information. In contrast to conventional GA applications, where problems are typically independent by nature (the optimisation of a function does not affect the optimisation of a different function ), this property of independence does not hold in the context of the background finder: the GA may identify a background involving a decay not listed in the PDG, and this newly discovered information could be relevant for other signals beyond the one currently under consideration. After the optimisation for a given signal is completed, the dictionary can be saved and reused for background identification in subsequent signals, avoiding redundant rediscovery of decays.
@@ 4 Environment
While the misidentification rate depends on the particle species and the experiment in question, we assume a global misidentification rate of 1% to validate our algorithmic approach, and write the misidentification penalty as , where denotes the number of misidentified particles.
@@ 5.1 GA performance
We conducted an initial experiment to evaluate the effectiveness of Genetic Algorithms (GAs) to identify the backgrounds associated with the training signals. Based on the toy model reward function, 94 backgrounds are considered relevant. To assess GA performance, defined as the number of relevant backgrounds identified, we performed multiple runs using varying population sizes and numbers of generations. The results, presented in Figure 8, demonstrate a clear upward trend in performance as either hyperparameter increases999Background processes may contain more final state particles than the signal if they involve non-reconstructed particles. For a more efficient search, instead of allowing the number of positive, negative and neutral final state particles in individuals to vary freely, we fixed these counts for the entire population. As a consequence, the process of finding the relevant backgrounds for a signal will in reality involve running multiple GA evolutions, each with a different configuration of these counts. For the signals in the considered training set, the number of evolutions is 3 or 4 depending on the case. The reported population sizes refer to those used in each of these evolutions.. This trend is expected: larger populations enhance genetic diversity, while more generations give the algorithm additional opportunities to explore the space. In Appendix B, we compute the GA space size for an example signal to be around . The results also indicate that, even with a total number of evaluations that is much smaller than the GA space size, the algorithm successfully identifies the majority of backgrounds relevant to the training signals.
@@ 5.4 Visualisation of learnt embeddings
6 Discussion
For the experiments in this work, we developed a toy scenario with 40 particles, using B-mesons as mother particles and experimentally well-established intermediate states, to provide realistic and sufficiently complex decay chains for evaluating the performance of the proposed algorithm. We conducted a first experiment to assess the performance of the GA, where results demonstrate that the GA can successfully identify most of the backgrounds for the training signals, despite using a much smaller number of evaluations than the size of the GA search space.
By employing the GA-discovered backgrounds as expert demonstrations, we trained and fine-tuned two agents, one following the PGSU approach and the other using PEG. The results indicate that both methods surpass expert performance and achieve perfect training performance101010In this context, perfect performance corresponds to correctly determining every background.. PGSU exhibits robust performance with slightly superior generalisation ability. In contrast, PEG achieves comparable results while providing faster initial performance improvement, as we only added a supervised term to the loss in Equation 5 for states along expert trajectories. This requires the agent to learn how to reach a state before it can receive supervision for that state. Consequently, PEG could be an interesting alternative for environments with very long episodes, where PGSU may result in a slower learning.
One of the most significant outcomes of this work is that it opens the door to a systematic strategy for accelerating and automating background determination in particle physics analyses. Traditionally, this task requires the physicist’s expert knowledge and manual inspection. However, by introducing an algorithmic approach based on ML, the procedure can be standardised, scaled to a wide range of analyses, and potentially integrated into large experimental pipelines.
The results presented in this work should be regarded as a proof of concept, demonstrating the feasibility of employing different AI and ML techniques to automate background determination. While the toy scenario employed here provides a controlled environment for testing the method, the encouraging performance of the proposed framework highlights its potential to be extended and adapted to real-world applications.
Nevertheless, several key challenges remain before such a framework can be deployed in realistic scenarios. First, in our experiments the training was performed using a small set of signal decays, and focused on a subset of 40 particles with constraints on the number of allowed intermediate resonances and misidentifications. While this setup provides a manageable and sufficiently complex environment, it limits the coverage of the space of possible background processes. In practice, training should be conducted with a much larger and more diverse set of signals to allow the agent to learn generalisable representations across the full decay space. Second, the toy model reward employed in our experiments was sufficient to validate the concept, but represents only a simplification of the complexities present in a realistic physics analysis. Future developments should explore more sophisticated, data-driven approaches to extract the reward, enabling the framework to handle realistic physics observables.
Furthermore, the GA component of our framework can be refined to further enhance its efficiency. In particular, the incorporation of additional conservation laws beyond charge conservation would constrain the search space and more effectively guide the exploration toward physically valid decay processes.
Taken together, these considerations highlight both the promise and the necessary next steps of this research direction. The methodology proposed here demonstrates that combining RL and GAs provides a powerful mechanism for addressing the highly-sparse-reward and large-search-space challenges inherent in background determination. With further refinement, including scaling to larger datasets, employing a more realistic reward function than the simplified one used here, and incorporating additional physics constraints into the algorithm, this approach could develop into a practical tool for high-energy physics analyses.
7 Conclusions
@@
The experiments conducted in this work demonstrate the effectiveness of the proposed algorithm. The GA efficiently explores and optimises the vast space of possible background processes, successfully identifying the majority of relevant backgrounds while requiring significantly fewer evaluations than the size of the full search space. By leveraging these GA-discovered backgrounds as expert demonstrations, both the PGSU and PEG techniques surpass expert-level performance, with PGSU achieving slightly superior generalisation ability, and PEG providing faster initial performance improvement. Furthermore, the transformer-based agent learns meaningful internal representations, with embeddings clustering according to particle composition and intermediate resonance structure. Together, these results validate the efficiency and learning capability of the proposed algorithm and highlight its potential to generalise beyond the training signals, providing a promising strategy for background determination in particle physics.
While the presented proof-of-concept experiments provide valuable insights, several limitations should be acknowledged. The training was conducted on a restricted set of particles, with constraints on the number of allowed intermediate resonances and misidentifications, and utilising a simplified reward model. This setup constrains the range of background processes considered, and simplifies complexities of real analyses. Nevertheless, the promising results highlight the framework’s potential to be expanded and applied to real-world scenarios, and demonstrate the effectiveness of the RL-GA synergy for addressing environments characterised by highly sparse rewards and a large trajectory space.
8 Conflicts of Interest
On behalf of all the authors, the corresponding author declares that there is no conflict of interest.
9 Data availability statement
10 Funding statement
The UZH group acknowledges support from the Swiss National Science Foundation (SNSF) under grants 209263 and 204238, and from Forschubngskredit under the UZH Candoc Grant no. FK-25-086. The Imperial College group acknowledges support from the STFC (United Kingdom) and The Leverhulme Trust (United Kingdom). The University of Bristol group acknowledges support from the STFC under grant ST/W000490/1.